On the topic of Forces and Motion, what is a velocity and acceleration vector?


1
Expert's answer
2021-01-19T07:07:57-0500

Let "\mathbf{r}" be the radius vector of a point. Then by definitio, the velocity vector is the derivative of "\mathbf{r}" with respect to time:


"\mathbf{v} \equiv \dfrac{d\mathbf{r}}{dt}"

The acceleration vector, by definition, is the derivative of "\mathbf{v}" with respect to time:


"\mathbf{a} \equiv \dfrac{d\mathbf{v}}{dt} = \dfrac{d^2\mathbf{r}}{dt^2}"